# A shared-revenue Bertrand game

### v1

@@ D.1. Shared-price equilibria $p_{r}=p=p_{s}$ in which the retailer fulfills demand
Condition I requires the retailer would not prefer to set a price $p^{\prime}<p$. Because $\pi_{r,r}$ is a concave quadratic function, it is increasing on $[0,p_{r^{*}}]$, so condition I holds if and only if $p\leq p_{r^{*}}$. Indeed, we derive an expression for $p_{r^{*}}$ by differentiating the quadratic $\pi_{r,r}$ and setting it equal to 00:
$\frac{\partial\pi_{r,r}}{\partial p_{r}}$ $=\frac{\partial}{\partial p_{r}}(p_{r}-c_{r})(1-p_{r})$
$=-2p_{r}+1+c_{r}\stackrel{{\scriptstyle\text{set}}}{{=}}0$
(36) $\implies p_{r^{*}}$ $=\frac{1+c_{r}}{2}$
Condition II is equivalent to saying that the retailer would not prefer to increase their price, effectively allowing the independent seller to fulfill demand at price $p$ instead of the retailer fulfilling demand at price $p$. Formally, we need
(37) $\pi_{r,s}(p)$ $\leq\pi_{r,r}(p)$
$\alpha p(1-p)$ $\leq(p-c_{r})(1-p)$
$\alpha p$ $\leq(p-c_{r})$
$p_{r_{\text{ind}}}=\frac{c_{r}}{1-\alpha}$ $\leq p$
Indeed, $p_{r_{\text{ind}}}$ is defined as the price at which the retailer’s payoff is equal no matter which seller fulfills demand, which is exactly equal to $\frac{c_{r}}{1-\alpha}$
$\pi_{r,r}(p_{r_{\text{ind}}})$ $=\pi_{r,s}(p_{r_{\text{ind}}})$
$(p_{r_{\text{ind}}}-c_{r})(1-p_{r_{\text{ind}}})$ $=\alpha p_{r_{\text{ind}}}(1-p_{r_{\text{ind}}})$
$-c_{r}$ $=\alpha p_{r_{\text{ind}}}-p_{r_{\text{ind}}}$
(38) $p_{r_{\text{ind}}}$ $=\frac{c_{r}}{1-\alpha}$
With this intuition, the simplification in (37) is almost true by definition. All that is required is to show the direction of the inequality, which is correct because the retailer’s payoff from the independent seller fulfilling demand at $p=0$ is zero, but is negative when it fulfills demand on its own at $p=0$.
Condition III posits that the independent seller would not prefer to undercut the retailer and fulfill demand themselves at some price $p^{\prime}<p$. However, recall that in an equilibrium where the retailer fulfills demand, the independent seller achieves a payoff of 00. Thus, we can rewrite condition III as:
@@
Because $\pi_{s,s}$ is a concave quadratic function, we have that
@@
Where $p_{s^{*}}$ is the maximum of $\pi_{s,s}$, for which we find a closed-form –
$\frac{\partial\pi_{s,s}}{\partial p_{s}}$ $=\frac{\partial}{\partial p_{s}}((1-\alpha)p_{s}-c_{s})(1-p_{s})$
$=-2(1-\alpha)p_{s}+(1-\alpha)+c_{s}\stackrel{{\scriptstyle\text{set}}}{{=}}0$
(40) $\implies p_{s^{*}}$ $=\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
@@
(41) $\pi_{s,s}(p)$ $\leq 0$
$((1-\alpha)p-c_{s})(1-p)$ $\leq 0$
$p$ $\leq\frac{c_{s}}{1-\alpha}=p_{s_{\text{ind}}}$
where we ignored the $1\leq p$ because $p\in[0,1]$. Indeed, we have that $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}$ because
$\pi_{s,s}(p_{s_{\text{ind}}})$ $=\pi_{s,r}(p_{s_{\text{ind}}})$
$((1-\alpha)p_{s_{\text{ind}}}-c_{s})(1-p_{s_{\text{ind}}})$ $=0$
(42) $p_{s_{\text{ind}}}$ $=\frac{c_{s}}{1-\alpha}$
@@ D.2. Shared-price equilibria $p_{r}=p=p_{s}$ in which the independent seller fulfills demand
Condition III requires that the independent seller would not prefer to fulfill demand at $p^{\prime}<p$. Because $\pi_{s,s}$ is a concave quadratic function maximized at $p_{s^{*}}$, condition III is equivalent to $p\leq p_{s^{*}}$.
Condition IV requires that the independent seller would not prefer that the retailer fulfill demand; it simplifies to
(44) $\pi_{s,r}(p)$ $\leq\pi_{s,s}(p)$
00 0 $\leq((1-\alpha)p-c_{s})(1-p)$
$p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}$ $\leq p$
@@
Because $\pi_{r,r}$ is a concave quadratic function, we have that
@@ D.2.1. Equilibria where $p\leq p_{r^{*}}$
Consider the first case of (45) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p)$. Condition I then simply becomes $\pi_{r,r}(p)\leq\pi_{r,s}(p)$, which we showed in (37) is equivalent to $p\leq p_{r_{\text{ind}}}$. Thus, we have found a continuum of equilibria $p\in\left[\max\{0,p_{s_{\text{ind}}}\},\min\{1,p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$, so long as the interval is nonempty. Note that $p_{s_{\text{ind}}}\geq 0$ and $p_{r^{*}}\leq 1$ always, so we can drop those.
@@
We showed in (43) that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always, and we have the following condition for $p_{s_{\text{ind}}}\leq p_{r^{*}}$
$p_{s_{\text{ind}}}$ $\leq p_{r^{*}}$
$\frac{c_{s}}{1-\alpha}$ $\leq\frac{1+c_{r}}{2}$
$2c_{s}$ $\leq(1-\alpha)(1+c_{r})$
(46) $\alpha$ $\leq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
Where $\alpha_{r^{*}}$ gets its name because $p_{r^{*}}\geq p_{s_{\text{ind}}}\iff\alpha\geq\alpha_{r^{*}}$ is necessary for $(p_{r^{*}},r)$ to be an equilibrium. Furthermore, we can see that $p_{s_{\text{ind}}}\leq p_{s^{*}}$ is equivalent to the following –
$p_{s_{\text{ind}}}$ $\leq p_{s^{*}}$
$\frac{c_{s}}{1-\alpha}$ $\leq\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
$2c_{s}$ $\leq 1-\alpha+c_{s}$
(47) $\alpha$ $\leq 1-c_{s}$
However, this is implied by $\alpha\leq\alpha_{r^{*}}$; we can see easily that $\alpha_{r^{*}}\leq 1-c_{s}$ because
(48) $\alpha_{r^{*}}=1-\frac{2c_{s}}{1+c_{r}}$ $\leq 1-c_{s}$
$c_{s}$ $\leq\frac{2c_{s}}{1+c_{r}}$
$c_{r}$ $\leq 1$
This intuitively makes sense, because as we argued in (9), the independent seller cannot attain positive payoff if $\alpha\geq 1-c_{s}$, so of course in this limit the equilibrium $(p_{r^{*}},r)$ is feasible. Thus the interval of prices is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
@@
We have the following condition for $p_{r_{\text{ind}}}\leq p_{r^{*}}$
$p_{r_{\text{ind}}}$ $\leq p_{r^{*}}$
$\frac{c_{r}}{1-\alpha}$ $\leq\frac{1+c_{r}}{2}$
$2c_{r}$ $\leq 1+c_{r}-\alpha-\alpha c_{r}$
$\alpha(1+c_{r})$ $\leq 1-c_{r}$
(49) $\alpha$ $\leq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
It will become clearer in Appendix D.2.2 where the $\alpha^{\dagger}_{r}$ notation comes from. Furthermore, $p_{r_{\text{ind}}}\leq p_{s^{*}}$ when
$p_{r_{\text{ind}}}$ $\leq p_{s^{*}}$
$\frac{c_{r}}{1-\alpha}$ $\leq\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
$2c_{r}$ $\leq 1-\alpha+c_{s}$
(50) $\alpha$ $\leq 1-2c_{r}+c_{s}=\alpha_{s^{*}}$
Note that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$ always because
(51) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $<1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
$1-c_{r}$ $<1+c_{r}-2c_{s}$
$c_{s}$ $<c_{r}$
We acknowledge that it is not yet clear when $\alpha_{s^{*}}\stackrel{{\scriptstyle?}}{{\leq}}\alpha^{\dagger}_{r}$ or $\alpha_{s^{*}}\stackrel{{\scriptstyle?}}{{\leq}}\alpha_{r^{*}}$; this is a question to which we will return shortly. However, we can conclude for now that $p_{r_{\text{ind}}}$ is the upper bound when $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$.
@@
We see that $p_{s^{*}}\leq p_{r^{*}}$ when
$p_{s^{*}}$ $\leq p_{r^{*}}$
$(1-\alpha)c_{r}$ $\geq c_{s}$
(52) $\alpha$ $\leq 1-\frac{c_{s}}{c_{r}}=\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
The notation $\alpha_{s^{*},r^{*}}$ comes from the implication of (52) between optimal prices for retailer and seller. Additionally, per (50), we have that $p_{s^{*}}\leq p_{r_{\text{ind}}}\iff\alpha\geq\alpha_{s^{*}}$, which is exactly the namesake of $\alpha_{s^{*}}$. Furthermore, we can see that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$
(53) $\alpha_{s^{*},r^{*}}=\frac{c_{r}-c_{s}}{c_{r}}=1-\frac{c_{s}}{c_{r}}$ $\leq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
This should once again be intuitive because clearly, $(p_{r^{*}},r)$ could not be an equilibrium if $p_{s^{*}}\leq p_{r^{*}}$, as the independent seller would prefer to undercut the retailer and sell at $p_{s^{*}}$ instead. Thus, we have that $p_{s^{*}}$ is the upper bound of the interval when $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, provided the interval is nonempty.
@@
Combining (52) and (49), we can immediately conclude that $p_{r^{*}}$ is the upper bound of the interval whenever $\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha_{r^{*}}\right]$.
@@
Explicitly simplifying the condition $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}$ is uninsightful, so instead we will break up the equilibria into two cases based on which direction of this inequality is true. However, we will prove one equivalence that will make our presentation easier throughout –
$\alpha_{s^{*}}=1-2c_{r}+c_{s}$ $\leq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
$1-c_{r}-(c_{r}-c_{s})$ $\leq\frac{1-c_{r}}{1+c_{r}}$
$1-\frac{c_{r}-c_{s}}{1-c_{r}}$ $\leq\frac{1}{1+c_{r}}$
$\frac{c_{r}}{1+c_{r}}$ $\leq\frac{c_{r}-c_{s}}{1-c_{r}}$
(54) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\leq\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
Note we showed in (51) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$, so as a corollary of (54), we know that $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}$ implies that $\alpha_{s^{*}}\leq\alpha_{r^{*}}$ as we wanted before. Finally, for ease of presentation, we will manipulate (54) to find an equivalent condition in terms of $c_{s}$ and $c_{s^{*}}$
$\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\leq 1-\frac{c_{s}}{c_{r}}=\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
$\frac{c_{s}}{c_{r}}$ $\leq\frac{2c_{r}}{1+c_{r}}$
(55) $c_{s}$ $\leq\frac{2c_{r}^{2}}{1+c_{r}}=c_{s^{*}}$
To summarize, we have so far found the following equilibria:
@@
(56) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha_{s^{*}}\\ \left[p_{s_{\text{ind}}},p_{s^{*}}\right]&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ \left[p_{s_{\text{ind}}},p_{r^{*}}\right]&\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\\ \end{cases}$
@@
(57) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha^{\dagger}_{r}\\ \left[p_{s_{\text{ind}}},p_{r^{*}}\right]&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\\ \end{cases}$
@@ D.2.2. Equilibria where $p\geq p_{r^{*}}$
Now, consider the second case of (45) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p_{r^{*}})$. The price at which $\pi_{r,s}(p)$ is equal to $\pi_{r,r}(p_{r^{*}})$ was defined in Section 2.2 to be $p^{\dagger}$, which is the solution to the following quadratic equation:
(58) $\pi_{r,r}(p_{r^{*}})$ $=\pi_{r,s}(p^{\dagger})$
$\frac{(1-c_{r})^{2}}{4}$ $=\alpha p^{\dagger}(1-p^{\dagger})$
$(p^{\dagger})^{2}-p^{\dagger}+\frac{(1-c_{r})^{2}}{4\alpha}$ $=0$
$p^{\dagger}$ $=\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$
Where we choose the larger solution, as the smaller solution satisfies the following inequalities
(59) $\frac{1-\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}<\frac{1}{2}<\frac{1+c_{r}}{2}=p_{r^{*}}$
@@
and therefore drops when we consider that we are handling the case of $p_{r^{*}}<p$. Note that, in order for $p^{\dagger}$ to be well defined, we need the discriminant to be positive which occurs precisely when $\alpha\geq(1-c_{r})^{2}$; we will return to this later in this subsection. Thus for an equilibrium in this case, we need that $p\in\left[\max\{0,p_{r^{*}},p_{s_{\text{ind}}}\},\min\{1,p^{\dagger},p_{s^{*}}\}\right]$. Note that $p^{\dagger}\leq 1$ and $p_{r^{*}},p_{s_{\text{ind}}}\geq 0$ always, so we can drop those conditions.
From (46), we have that $p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}$. We won’t explicitly simplify the condition $p_{s^{*}}\leq p^{\dagger}$ here; we delay that until (76). For now we will simply remark that $\pi_{r,s}(p)$ is maximized at $p=\frac{1}{2}$, which follows because
$\frac{\partial\pi_{r,s}}{\partial p}$ $=\frac{\partial}{\partial p}\left(\alpha p(1-p)\right)$
$=\alpha(1-2p)\stackrel{{\scriptstyle\text{set}}}{{=}}0\iff p=\frac{1}{2}$
So, because $\frac{1}{2}\leq p_{s^{*}},p^{\dagger}$ and $\pi_{r,s}$ is a concave quadratic, we have that
(60) $p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})\leq\pi_{r,s}(p_{s^{*}})$
We’ll check to make sure that the interval of prices is nonempty in each of the possible cases:
Interval is $[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}]$.
Per (47), we have that $p_{s_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq 1-c_{s}$. However, we have not found the conditions under which $p_{s_{\text{ind}}}\leq p^{\dagger}$. By the same argument as in (60), we have the following equivalences:
$p_{s_{\text{ind}}}$ $\leq p^{\dagger}$
$\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$ $\leq\pi_{r,s}(p_{s_{\text{ind}}})$
(61) $\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha c_{s}}{1-\alpha}\left(1-\frac{c_{s}}{1-\alpha}\right)$
Now, recall that
$p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}\geq\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}\geq(1-c_{r})^{2}$
which implies that $p^{\dagger}$ is well defined and that $\pi_{r,s}(\frac{1}{2})\geq\pi_{r,r}(p_{r^{*}})$. We see upon substituting the endpoint $\alpha=\alpha^{\dagger}_{r}=1-\frac{2c_{s}}{1+c_{r}}$ into (61) that
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha^{\dagger}_{r}c_{s}}{1-\alpha^{\dagger}_{r}}\left(1-\frac{c_{s}}{1-\alpha^{\dagger}_{r}}\right)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\left(\frac{1-2c_{s}+c_{r}}{1+c_{r}}\right)\frac{c_{s}}{\frac{2c_{s}}{1+c_{r}}}\left(1-\frac{c_{s}}{\frac{2c_{s}}{1+c_{r}}}\right)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\left(\frac{1-c_{r}}{2}\right)\left(\frac{1-2c_{s}+c_{r}}{2}\right)$
$1-c_{r}$ $\leq 1+c_{r}-2c_{s}$
$c_{s}$ $\leq c_{r}$
Furthermore, we show in (85) and (86) that $\pi_{r,s}(p_{s_{\text{ind}}},\alpha)$ has a unique maximum at $\alpha=\frac{1-c_{s}}{1+c_{s}}<\alpha_{r^{*}}$, which is sufficient to imply that there exists some $\alpha^{\dagger}_{s}\in\left[\alpha_{r^{*}},1-c_{s}\right]$ such that $\alpha\leq\alpha^{\dagger}_{s}\iff p_{s_{\text{ind}}}\leq p^{\dagger}$. For completeness, explicitly solving (61) yields the closed-form
(62) $\alpha^{\dagger}_{s}\equiv\frac{c_{s}(1-c_{s})+\frac{(1-c_{r})^{2}}{2}+c_{s}\sqrt{(1-c_{s})^{2}-(1-c_{r})^{2}}}{2c_{s}+\frac{(1-c_{r})^{2}}{2}}\in\left[\alpha_{r^{*}},1-c_{s}\right]$
Therefore, the interval is feasible as long as $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$.
Interval is $[p_{r^{*}},p_{s^{*}}]$.
Per (52), we have that $p_{r^{*}}\leq p_{s^{*}}\iff\alpha\geq\alpha_{s^{*},r^{*}}$, so $[p_{r^{*}},p_{s^{*}}]$ is nonempty whenever $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha_{r^{*}}\right]$. Note we showed in (53) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$ always.
Interval is $[p_{r^{*}},p^{\dagger}]$.
Interestingly, we find the following condition for $p_{r^{*}}\leq p^{\dagger}$:
$p^{\dagger}$ $\geq p_{r^{*}}$
$\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$ $\geq\frac{1+c_{r}}{2}$
$\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}$ $\geq c_{r}$
$\frac{(1-c_{r})^{2}}{\alpha}$ $\leq 1-c_{r}^{2}$
(63) $\alpha$ $\geq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
@@
Thus, $[p_{r^{*}},p^{\dagger}]$ is nonempty so long as $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{r^{*}}\right]$. Note we showed in (51) that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$ always.
Reconciling equilibria with case (i) where $c_{s}\leq c_{s^{*}}$.

### v2

@@ D.1. Shared-price equilibria $p_{r}=p=p_{s}$ in which the retailer fulfills demand
Condition I requires the retailer would not prefer to set a price $p^{\prime}<p$. Because $\pi_{r,r}$ is concave, it is increasing on $0,p_{r^{*}}]$, so condition I holds if and only if $p\leq p_{r^{*}}$.
Condition II is equivalent to saying that the retailer would not prefer to increase their price, effectively allowing the independent seller to fulfill demand at price $p$ instead of the retailer fulfilling demand at price $p$, which by construction is equivalent to $p\geq p_{r_{\text{ind}}}$.
Condition III posits that the independent seller would not prefer to undercut the retailer and fulfill demand themselves at some price $p^{\prime}<p$. However, recall that in an equilibrium where the retailer fulfills demand, the independent seller achieves a payoff of 0. Thus, we can rewrite condition III as:
@@
Because $\pi_{s,s}$ is a concave unimodal function uniquely maximized at $p_{s^{*}}$, we have that
@@
(45) $\pi_{s,s}(p)\leq 0=\pi_{s,r}(p)\iff p\leq\frac{c_{s}}{1-\alpha}=p_{s_{\text{ind}}}$
where we ignored the $1\leq p$ because $p\in[0,1]$.
@@ D.2. Shared-price equilibria $p_{r}=p=p_{s}$ in which the independent seller fulfills demand
Condition III requires that the independent seller would not prefer to fulfill demand at $p^{\prime}<p$. Because $\pi_{s,s}$ is a concave unimodal function maximized at $p_{s^{*}}$, condition III is equivalent to $p\leq p_{s^{*}}$.
Condition IV requires that the independent seller would not prefer that the retailer fulfill demand, in other words $p\geq p_{s_{\text{ind}}}$.
@@
Because $\pi_{r,r}$ is a concave unimodal function, we have that
@@ D.2.1. Equilibria where $p\leq p_{r^{*}}$
Consider the first case of (47) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p)$. Condition I then simply becomes $\pi_{r,r}(p)\leq\pi_{r,s}(p)$, which by construction is equivalent to $p\leq p_{r_{\text{ind}}}$. Thus, we have found a continuum of equilibria $p\in\left[\max\{0,p_{s_{\text{ind}}}\},\min\{1,p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$, so long as the interval is nonempty. Note that $p_{s_{\text{ind}}}\geq 0$ and $p_{r^{*}}\leq 1$ always, so we can drop those.
@@
We showed in (46) that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always, and $p_{s_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{r^{*}}$. Thus the interval of prices is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
@@
Recall that $p_{r_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}$, and $p_{r_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq\alpha_{s^{*}}$. Thus, we need $\alpha\leq\min\{\alpha^{\dagger}_{r},\alpha_{s^{*}}\}$ in order for $p_{r_{\text{ind}}}$ to be the upper bound of the interval. Note also that $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}\iff c_{s}\leq c_{s^{*}}$.
@@
We have that $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$ and $p_{s^{*}}\leq p_{s_{\text{ind}}}\iff\alpha\geq\alpha_{s^{*}}$, thus this case occurs if $\alpha\in[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}]$. Also, recall that $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}\iff c_{s}\leq c_{s^{*}}$.
@@
Since we need both $\alpha>\alpha_{s^{*},r^{*}}$ and $\alpha>\alpha^{\dagger}_{r}$ in order for $p_{r^{*}}$ to b the upper bound, this occurs whenever $\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha_{r^{*}}\right]$.
@@
We have so far found the following equilibria:
@@
[…]
@@
[…]
@@ D.2.2. Equilibria where $p\geq p_{r^{*}}$
Now, consider the second case of (47) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p_{r^{*}})$. The largest price for which $\pi_{r,s}(p)$ is equal to $\pi_{r,r}(p_{r^{*}})$ was defined in Section 2.2 to be $p^{\dagger}$. Thus for an equilibrium in this case, we need that $p\in\left[\max\{0,p_{r^{*}},p_{s_{\text{ind}}}\},\min\{1,p^{\dagger},p_{s^{*}}\}\right]$. Note that $p^{\dagger}\leq 1$ and $p_{r^{*}},p_{s_{\text{ind}}}\geq 0$ always, so we can drop those conditions.
Recall that $p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}$. We won’t explicitly simplify the condition $p_{s^{*}}\leq p^{\dagger}$, but note in light of Lemma 2 we have
$p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})\leq\pi_{r,s}(p_{s^{*}})$
Interval is $p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}]$.
@@
Recall that $p_{s_{\text{ind}}}\leq p_{s^{*}}$ always, and $\alpha\in[\alpha_{r^{*}},\alpha^{\dagger}_{s}]\implies p_{s_{\text{ind}}}\leq p^{\dagger}$ per Corollary 5. Therefore, the interval is feasible as long as $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$, which we showed above is always nonempty.
Interval is $p_{r^{*}},p_{s^{*}}]$.
Recall that $p_{r^{*}}\leq p_{s^{*}}\iff\alpha\geq\alpha_{s^{*},r^{*}}$, so $p_{r^{*}},p_{s^{*}}]$ is nonempty whenever $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha_{r^{*}}\right]$. Note we showed in (40) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$ always.
Interval is $p_{r^{*}},p^{\dagger}]$.
We have $p_{r^{*}},p^{\dagger}]$ is nonempty so long as $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{r^{*}}\right]$. Note we showed in (41) that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$ always.
Equilibria when $c_{s}\geq c_{s^{*}}$.